1. Use the following information to answer questions Lite Products, Inc. is considering investing in either of two competing projects that will allow the firm engineering department narrowed the alternatives down to two- Status Quo (SQ) and High Tech estimates of the cash flows for SQ and HT over the relevant six-year time horizon. The firm has an 11 percent required return and views these projects as equally risky. Project SQ     Project HT Initial Outflow (CFo)   $670,000  $940,000 Year (t)                             Cash Inflows (CFt)   1    $250,000  $170,000   2   200,000  180,000   3    170,000  200,000 4    150,000  250,000 5    130,000  300,000 6    130,000  550,000 A. Based on the information given above, calculate the net present value (NPV) for each asset respectively. (Round to the nearest four decimal points or use table value).B. Which asset is best, using the NPV? 


1
Expert's answer
2022-01-02T18:20:24-0500

Solution:

AStatus Quo (SQ) NPV:

NPV =  "\sum \frac{CF}{(1 + r)^{t} }" – Initial Investment (CFo)


NPV = "(\frac{250,000}{(1 + 0.11)^{1} } + \frac{200,000}{(1 + 0.11)^{2} } + \frac{170,000}{(1 + 0.11)^{3} } + \frac{150,000}{(1 + 0.11)^{4} } + \frac{130,000}{(1 + 0.11)^{5} }+\frac{130,000}{(1 + 0.11)^{6} } ) - 670,000"

= (225,225.23 + 162,324.49 + 124,302.53 + 98,809.65 + 77,148.67 + 69,503.31) – 670,000

= 757,313.88 – 670,000 = 87,313.88

NPV = 87,313.88

 

High Tech NPV:

NPV = "(\frac{170,000}{(1 + 0.11)^{1} } + \frac{180,000}{(1 + 0.11)^{2} } + \frac{200,000}{(1 + 0.11)^{3} } + \frac{250,000}{(1 + 0.11)^{4} } + \frac{300,000}{(1 + 0.11)^{5} }+\frac{550,000}{(1 + 0.11)^{6} } ) - 940,000"

= (153,153.15 + 137,975.81 + 146,238.23 + 164,682.74 + 178,035.40 + 294,052.55) – 940,000

= 1,074,137.86 – 940,000 = 134,137.86

NPV = 134,137.86

 

B). The best asset using NPV is High Tech.

Both projects are acceptable since they have a positive NPV.

However, High Tech has a much larger positive NPV than Status Quo